OpenStudy (anonymous):

divide the polynomial by the monomial (27x^9-72x^8+5x^7-45x^6)/(-9x^6)

OpenStudy (anonymous):

You divide each term by -9x^6 -3x^3 + 8x^2 -5/9x + 5

OpenStudy (anonymous):

Remember that \[\Large \frac{{{x^n}}}{{{x^m}}} = {x^{n - m}}\]

OpenStudy (anonymous):

what about the -45x^6?

OpenStudy (anonymous):

That becomes 5, since -45/-9 = 5 and x^6/x^6 = 1. I got: \[\Large - 3{x^3} + 8{x^2} - \frac{5}{9}x + 5\]

OpenStudy (anonymous):

... which is actually exactly the same as blexting's, I just read his incorrectly.